i need help on this problem fast, i need the answer in the next hour...please! this was the only problem that i cant solve. the price,p, of oil varies inversely as the supply, s. an opec nation sells oil for $19.50 per barrel when its daily production level is 4 million barrels. at what price will it sell if the daily production level is decreased by 3 million barrels?
Answer by stanbon(75887)

If they vary inversely p=k/s where "k" has to be determined.
If price =18.5 when s=4000000 then 19.5=k/4000000 and k=78000000
So your equation is p=78000000/s
If te amount of oil is decreases by 3000000 the supply will
be 1000000 and s=78000000/1000000=$78
Cheers,
Stan H.